Q. 1. The joint pdf of random variables ‘X’ and ‘Y’ is

 

Find

a. The marginal pdfs, fX(x) and fY(y).

b. The conditional pdfs, fX/Y(x/y) and fY/X(y/x)

c. The E(X/Y=1)

d. Are ‘X’ and ‘Y’ statistically independent?



1
Expert's answer
2021-05-11T06:04:26-0400

(a) "f_x(x)=\int f_{x,y}(x,y)dy=\int_x^2\frac{1}{2}dy=\{ \frac{2-x}{2},\ \ 0\le x\le 2 \ and \ 0,\ \ else"


"f_y(y)=\int f_{x,y}(x,y)dx=\int_0^y\frac{1}{2}dx=\{\frac{y}{2},0\le y\le 2\ \ and \ \ 0, \ else"



(b) "f_{\frac{X}{Y}}(X/Y)=\dfrac{f_{x,y}(x,y)}{f_y(y)}=\dfrac{1/2}{y/x}=\{\frac{1}{y}, \ \ 0\le x\le y\ \ and\ \ 0,\ \ \ else"


"f_{Y/X}(Y/X)=\dfrac{f_{x,y}(x,y)}{f_x(x)}=\dfrac{1/2}{(2-x)/2}=\{\frac{1}{2-x},\ \ x\le y\le 2\ \ \ and \ 0,\ \ else"


(c) "E(X/Y=1)=\int_0^1f_{X/Y}(x/1)dx=\int_0^11dx=1"


(d) "f_{X/Y}(x/y)\neq f_x(x)\\"

"\Rightarrow X, Y" are not statistically independent.